OpenStudy (anonymous):

Find the discriminant and use it to determine if the solution has one real, two real, or two imaginary solutions. x^2+4x=1=0 SHOW WORK Find the discriminant and use it to determine if the solution has one real, two real, or two imaginary solutions. x^2+4x=1=0 SHOW WORK @Mathematics

jimthompson5910 (jim_thompson5910):

From x^2+4x+1, we see that a = 1, b = 4, and c = 1. Plug these into the discriminant formula D=b^2-4ac to get D=b^2-4ac D=(4)^2-4(1)(1) D=16-4(1)(1) D=16-4 D=12 So the discriminant is D=12 Since the discriminant is greater than zero, this means that there are two real solutions.